OpenStudy (anonymous):

A cat stands on a plank of wood floating in the water. The mass of the cat is 3.5 kg and the mass of the wood is 5.0 kg. If the cat walks 1.0 m along the plank, how far does it move in relation to the water?

OpenStudy (anonymous):

The center of gravity of the cat-plank system does not move. cat mass * distance moved in space right = plank mass *distance moved in space left. 3.5 c = 5 p c+p = 1 meter so p = 1-c 3.5 c = 5 (1-c) 8.5 c = 5 c = .588 meter